### v1

# Structural modulation, physical properties, and electronic band structure of the kagome metal UCr6Ge6

Abstract
The chemical flexibility of the stoichiometry, where an -block element is intercalated in the CoSn structure type, allows for the tuning of flatbands associated with kagome lattices to the Fermi level and for emergent phenomena due to interactions between the - and -electron lattices. Yet, 5 members of the “166” compounds are underrepresented compared with 4 members. Here, we report single-crystal growth of UCr6Ge6, which crystallizes in a monoclinically distorted Y0.5Co3Ge3-type structure. The real-space character of the modulation, which is unique within the family, is approximated by a 312 supercell of the average monoclinic cell. The compound has kagome-lattice flatbands near the Fermi level and a moderately enhanced electronic heat capacity, as evidenced by its low-temperature Sommerfeld coefficient ( mJ mol-1 K-2) paired with band structure calculations. The small, isotropic magnetization and featureless resistivity of UCr6Ge6 suggest itinerant uranium 5 electrons and Pauli paramagnetism. The isotropic magnetic behavior of the uranium 5 electrons starkly contrasts with localized behavior in other uranium 166 compounds, highlighting the high tunability of the magnetic ground state across the material family.
I Introduction
Materials with the stoichiometry (=Sc, Y, -block element; =transition metal; =Ga, Si, Ge, Sn) contain a kagome transition metal lattice intercalated by an element. The interplay of these sublattices often produces emergent phenomena. For example, in (“166”) materials with small elements relative to the cages surrounding them, charge density wave instabilities can occur, as in ScV6Sn6 and LuNb6Sn6 [arachchige2022charge, tuniz2023dynamics, pokharel2023frustrated, kim2023infrared, meier2023tiny, hu2024phonon, yi2024tuning, ortiz2024stability, meier2025pressure]. In systems with magnetic or elements, anisotropic magnetic behavior may produce complex magnetic phase diagrams involving incommensurate phases or noncollinear spin structures that produce large topological Hall effects [clatterbuck1999magnetic, kimura2006high, riberolles2024new, ghimire2020competing, l2025high, dhakal2021anisotropically, riedel2025magnetic]. Moreover, the extensive chemical tunability of the 166 family offers a platform for controlling electronic and structural behavior, such as Fermi level () density of states (DOS) enhancement due to kagome-lattice flatbands or reduced dimensionality due to well-separated kagome layers. However, despite numerous reported 166 materials, most show no more than a moderate enhancement in the DOS at , and only a handful have reported Sommerfeld coefficients exceeding 60 mJ mol-1 K-2 [pokharel2021electronic, ishii2013ycr6ge6, avila2005direct, lyu2024anomalous, shi2025weak, pokharel2022highly, guo2023triangular, lou2025strongly], all in the V6Sn6, Cr6Ge6, Mn6Sn6, Fe6Ge6, and Fe6Sn6 subfamilies.
Reports of actinide 166 materials are sparse compared with those for lanthanide 166s. Therefore, actinides’ 5 electrons offer a new route to tune electronic behavior associated with kagome-lattice flatbands. Reported actinide 166s include UCo6Ge6 [buchholz1981intermetallische], UFe6Ge6 [gonccalves1994ufe6ge6, waerenborgh2005crystal], UV6Sn6 [thomas2025uv6sn6, patino2025incom], ThV6Sn6 [thomas2025uv6sn6, xiao2024preparation], UNb6Sn6 [riedel2025magnetic], and ThNb6Sn6 [riedel2025magnetic]. Recent work on UV6Sn6 and UNb6Sn6 has shown that the -electron flatband is closer to and less dispersive for UV6Sn6 [riedel2025magnetic, thomas2025uv6sn6], indicating that chemical substitutions may significantly tune the energy level and shape of the kagome-lattice flatband. Here, we report a new uranium member of the Cr6Ge6 subfamily [mulder1993Gd, schobinger1997atomic, schobinger1997ferrimagnetism, konyk2020electrical, yang2024crystal, romaka2024structure, lee2025coexisting], which includes YCr6Ge6 and YbCr6Ge6, two compounds with kagome-lattice flatbands close to in the calculated band structures [ishii2013ycr6ge6, lee2025coexisting]. We report on the single-crystal synthesis and physical properties of UCr6Ge6, a compound with unique magnetic behavior and the largest electronic heat capacity among the reported actinide 166s.
II Methods
II.1 Experimental
Single crystals of UCr6Ge6 were grown from tin flux. Depleted uranium (99.99%), chromium (Thermo Scientific, 99.997%), germanium (Thermo Scientific, 99.9999+%), and tin (Thermo Scientific, 99.999%) pieces were placed in an alumina crucible in a 1:6:18:100 or 1.5:6:18:100 molar ratio. Samples were heated to 1100C at 100C/h, homogenized at 1100C for 72 h, and then slow cooled at 2C/h to 800C, where they were centrifuged to remove the tin flux. For the 1.5:6:18:100 ratio, hexagonal plate crystals were recovered. For the 1:6:18:100 ratio, hexagonal plates as well as hexagonal rods were recovered. Additionally, needle-shaped single crystals of LuCr6Ge6 were grown in tin flux using a 1:6:6:20 (Lu:Cr:Ge:Sn) molar ratio and were spun at 500∘C, with LuCrxGe2 crystals forming as a by-product (CeNiSi2-type, ).
The crystal structure of UCr6Ge6 was determined using data collected with a Bruker D8 VENTURE KAPPA single crystal X-ray diffractometer with an IS 3.0 microfocus source ( 0.71073 Å), a HELIOS optics monocromator, and a PHOTON II CPAD detector. All data were integrated with SAINT V8.42 [saint2025bruker], yielding 5733 reflections, of which 311 where independent (average redundancy 18.43) and 100% were greater than 2(F2). A Multi-Scan absorption correction using SADABS-2016/2 [krause2015comparison] was applied. The structure was solved by Intrinsic Phasing methods with SHELXT-2018/2 [sheldrick2015shelxt] and refined by full-matrix least-squares methods against F2 using SHELXL-2019/2 [sheldrick2015crystal]. All atoms were refined with anisotropic displacement parameters. Single crystal diffraction of LuCr6Ge6 followed the same general procedure. Crystallographic data for the structures reported in this paper have been deposited with the Cambridge Crystallographic Data Centre [groom2016cambridge], and CIFs are provided in the Supplemental Material [supplement].
To collect heat capacity data, the two-tau thermal relaxation method was used with a Quantum Design Physical Property Measurement System (PPMS). A 3He fridge attachment was used for data collected below 1.8 K. For resistivity measurements below 400 K and up to 9 T, platinum wires in a four-point configuration were attached to UCr6Ge6 crystals by spot welding. The crystal orientation was confirmed with a Photonic Science Laue diffractometer. The kagome plane was found to be parallel to the face of the plate-like crystals and perpendicular to the long axis of the rod-like crystals. Data collected under a magnetic field were averaged at positive and negative fields [( + )/2] to account for any small offset between the voltage leads. Energy dispersive spectroscopy (EDS) maps were collected with a ThermoFisher Apreo 2 S scanning electron microscope and were processed with the Oxford program AZtec.
II.2 Theoretical
Density functional theory (DFT) calculations were carried out using the pseudopotential projector-augmented wave method [kresse1999ultrasoft] implemented in the Vienna Ab initio Simulation Package (VASP) [kresse1993ab, kresse1996efficient]. An energy cutoff of 320 eV was used for the plane-wave basis set. Exchange-correlation effects were treated using the Perdew-Burke-Ernzerhof (PBE) generalized gradient approximation density functional [perdew1996generalized]. An 11117 -centered k-point mesh was used to sample the Brillouin zone. Spin-orbit coupling effects were included self-consistently. The DFT+ calculations utilized an effective Hubbard of 6 eV on the uranium 5 or the lutetium 4 states [dudarev1998electron]. The UCr6Ge6 unit cell was constructed using a simplified cell with lattice parameters , , where and are the lattice parameters of the average monoclinic cell discussed in III.1. A total energy tolerance of 10-6 eV was used to determine the self-consistent charge density.
III Results and Discussion
III.1 Crystal structure and composition
UCr6Ge6 crystallizes in a modulated structure with the average unit cell having space group symmetry. The lattice parameters of the average cell are Å, Å, Å, and (Fig. 1), and the corresponding modulation vector is , within refinement error of a commensurate vector . Additional refinement details are provided in the Supplemental Material [supplement]. For other stuffed CoSn materials, including UV6Sn6 [patino2025incom, thomas2025uv6sn6], modulated intergrowths of the ScFe6Ga6 () and ScFe6Ge6 () structure types are common [fredrickson2008origins]. However, the modulation behavior of UCr6Ge6 is distinct from that of other 166s.
The average cell is better described by the chemical formula U0.5Cr3Ge3 (analogous to the Y0.5Co3Ge3 structure [buchholz1981intermetallische]), but for simplicity in comparisons to other 166 compounds, we will use UCr6Ge6 as the formula unit (f.u.) for normalizations. A 1/2 component for UCr6Ge6 can be viewed as returning the unit cell length along to that of the full 1-6-6 cell, rather than that of the half 0.5-3-3 cell, as is also the case for UV6Sn6 [thomas2025uv6sn6]. A 2/3 modulation would triple the real-space axis. The 2/3 (in-plane) component is distinct from that of other 166 intergrowths, for example TbFeGe3.5Ga2.5 () [venturini2001crystallographic, venturini2006filling, fredrickson2008origins], where the real-space propagation of the modulation in the kagome plane has a component perpendicular to that of UCr6Ge6, i.e. along .
To visualize the unit cell modulation, the modulation reflections [circled in white in Fig. 1(c)] are refined with the average-cell reflections [circled in pink in Fig. 1(c)]. The combined reflection set is reasonably modeled by a cell with space-group symmetry (, ). Since is a supergroup of [ivantchev2000subgroupgraph], the refined hexagonal cell can be transformed into a monoclinic cell with nominal symmetry. The lattice vectors of the resulting monoclinic cell correspond to a supercell of the average monoclinic cell, and the periodicity matches a modulation (Fig. 2), indicating it is an approximate representation of the modulation in real space. Additional details on the relationship between the cell and the cells are provided in the Supplemental Material [supplement].
The supercell contains two unique, disordered U-Ge channels running along the axis. One contains half-occupied U and Ge sites (gray polyhedra), as in the Y0.5Co3Ge3-type cell, while the other is split between primary [87.7(6)%] and secondary [12.3(6)%] U and Ge sites (blue and orange polyhedra), matching the SmMn6Sn6-type disorder [malaman1997magnetic] observed in UNb6Sn6 [riedel2025magnetic]. As established for the SmMn6Sn6-type structure, the primary and secondary U/Ge sites cannot be occupied in neighboring cages along in the actual crystal; otherwise, bond distances are unreasonable because of overlap between the positions of the top and bottom Ge vertices in neighboring cages. Further analysis may require 3+1 superspace group symmetry treatments, which are beyond the scope of this work.
To confirm the 1:6:6 stoichiometry is maintained, EDS data was collected on the polished surface of a crystal (Fig. 3). The surface shows a uniform distribution of uranium, chromium, and germanium, giving an average U-Cr-Ge composition of 7.7(3)-46(2)-47(1) at.% over maps of four regions, which matches the expected 7.692-46.154-46.154 at.% ratio. The pictured region also contains a streak of residual tin flux not fully removed by polishing. The isolation of tin to the streak demonstrates the lack of tin incorporation in the bulk.
Structural analysis was also performed for the reference compound LuCr6Ge6. Single crystal diffraction did not show evidence of the site disorder reported for polycrystalline samples in Refs. [konyk2020electrical, romaka2024structure]. Instead, the data indicate a fully ordered cell. The discrepancy may be due to the lower effective quenching temperature in our work (500∘C) than for the arc-melted samples in the previous works. Additional refinement details are in the Supplemental Material [supplement].
III.2 Magnetic properties
The magnetic susceptibility, , of UCr6Ge6 is small and linear from 350 K to 30 K, where an onset indicative of magnetic order appears [Fig. 4(a)]. Magnetically ordered Cr6Ge6 materials are ferrimagnets with and Cr magnetic sublattices, Cr moment magnitudes around 0.5 , and moments reduced from the free-ion values [schobinger1997atomic, schobinger1997ferrimagnetism]. For the magnetically ordered lanthanide systems, a more even split in lanthanide occupation of the and lattice sites correlates with a higher magnetic ordering temperature [schobinger1997ferrimagnetism]. UCr6Ge6 represents the maximum limit of site disorder in that framework, where the and lattice sites are equally occupied, leading to a halving of the unit cell (ignoring the monoclinic modulation). A transition temperature higher than that of TbCr6Ge6 ( K [schobinger1997atomic]) is, therefore, reasonable. However, the magnetic susceptibility of UCr6Ge6 does not show Curie-Weiss behavior in the paramagnetic region. Though a fit to the relationship gives an effective moment of 8.84 /f.u. () or 9.49 /f.u. (), which superficially suggests magnetic chromium atoms similar to other Cr6Ge6 materials [brabers1994magnetic, yang2024crystal, romaka2024structure], the fits result in unrealistic Curie-Weiss temperatures () near -2000 K (Fig. S2 [supplement]). The shape and magnitude of , therefore, suggest that the 30 K transition is due to an impurity, and UCr6Ge6 is a Pauli paramagnet. In contrast, UV6Sn6 and UNb6Sn6, where the transition metals do not carry a magnetic moment, exhibit localized 5 electron magnetism [patino2025incom, thomas2025uv6sn6, riedel2025magnetic].
A review of possible Cr–Ge and U–Ge binary impurity phases finds no compounds with a transition near 30 K [zagryazhskii1968magnetic, kolenda1980esca, sato1983magnetic, sato1988electrical, onuki1992magnetic, troc2002magnetotransport, ghimire2012complex, ji2025small], and there are no reported U–Cr binary compounds [venkatraman1985CrU]. A possible ternary impurity phase is U3CrGe5, which has a reported transition near 25-30 K, though that transition may also be extrinsic [moussa2019overview]. Assuming the magnetic susceptibility jump in Ref. [moussa2019overview] and the feature in Fig. 4a are from U3CrGe5, the estimated impurity fraction is wt% of the sample. Moreover, the magnitude of the jump varied significantly between UCr6Ge6 samples (up to 4 larger than in Fig. 4).
If one assumes that the U electrons in UCr6Ge6 are itinerant and their Pauli paramagnetism contributes to the magnetic susceptibility, the estimated Pauli susceptibility is emu/mol, given a Wilson ratio [ =() ] of 1 and the measured Sommerfeld coefficient mJ/mol-K2 discussed below. This is 3.5 smaller than the measured value. If this interpretation is correct, then either there may be U ferromagnetic correlations, or a paramagnetic contribution from the Cr -electrons may dominate the magnetic susceptibility.
An alternative explanation for the small magnetic susceptibility of UCr6Ge6 is that the crystalline electric field (CEF) splits a ( valence) multiplet into 9 singlets, with a small van Vleck contribution to due to a large splitting between the nonmagnetic singlet ground state to another singlet first-excited state (perhaps with additional contributions to from the Cr -electrons). The Cr–Ge cages of UCr6Ge6 are smaller than the V–Sn and Nb–Sn cages in UV6Sn6 and UNb6Sn6. The uranium atoms that fill the Cr–Ge cages may, therefore, be more stable in a 5 valence configuration. The large number of symmetry-allowed crystal field parameters for the monoclinic structure means that crystal field fits to and are underconstrained. Therefore, further measurements, e.g., resonant inelastic x-ray scattering, are necessary to understand the valence state of the U electrons in UCr6Ge6.
The magnetization, , of UCr6Ge6 was measured at 1.8 K with the applied magnetic field parallel or perpendicular to the kagome lattice plane [Fig. 4(b)]. There is no appreciable anisotropy, and the moment size per formula unit is only 0.052 at 6.5 T. Other Cr6Ge6 compounds approach saturation near 7 T with magnetization around 1–10 /f.u. [brabers1994magnetic, yang2024crystal]. A small curvature below 0.5 T and a remnant magnetization of order 10-3 /f.u. were observed in some crystals of UCr6Ge6 where the impurity transition magnitude in was also larger, providing further evidence of a small ferromagnetic impurity leading to the 30 K feature. As with , the small and isotropic magnetization suggests itinerant uranium 5 electrons and paramagnetism. Similarly small values of the magnetization are observed for several Fe6Ge6 compounds with nonmagnetic elements, but in each, Fe orders well above room temperature [mazet2001macroscopic]. No decrease in with decreasing temperature is observed for UCr6Ge6, suggesting there is not an analogous high-temperature ordering transition of the chromium sublattice.
III.3 Heat capacity and band structure
The heat capacity of UCr6Ge6 plotted as (Fig. 5), is mostly featureless but exhibits two bumps near 30 K attributed to an impurity phase (Fig. S3 [supplement]). A fit of between K yields mJ mol-1 K-2 and mJ mol-1 K-4. The large Sommerfeld coefficient of UCr6Ge6 is comparable to that reported for LuFe6Ge6, YbFe6Ge6, and Lu1-xFe6Sn6 ( mJ mol-1 K-2 [avila2005direct, lyu2024anomalous, shi2025weak]), though smaller than that of YbV6Sn6 ( mJ mol-1 K-2 [guo2023triangular]) and YbMn6Sn6 ( mJ mol-1 K-2 [lou2025strongly]). Additionally, the Debye temperature of UCr6Ge6 is determined to be =444(4) K, where is the number of atoms per formula unit (13) and is the gas constant. is consistent with the larger values (calculated with DFT) for Cr6Ge6 compounds with lighter elements [romaka2024structure]. As a comparison, the heat capacity of LuCr6Ge6 was also collected (Fig. S4(a) [supplement]). The fit Sommerfeld coefficient and Debye temperature were mJ mol-1 K-2 and K for LuCr6Ge6. The Sommerfeld coefficient of LuCr6Ge6 is similar to that of YCr6Ge6 (66 mJ mol-1 K-2), which has electronic bands near associated with the chromium kagome lattice [ishii2013ycr6ge6]. The similarity suggests that the larger electronic heat capacity of UCr6Ge6 may be driven by band structure features associated with not only chromium but also additional uranium 5 states that increase the DOS near .
To assess whether the electronic heat capacity is influenced by chromium kagome flatbands and uranium 5 states, the electronic band structure was calculated for an ordered UCr6Ge6 cell with uranium 5 electrons in the valence [Fig. 6(a)] and for LuCr6Ge6 [Fig. 6(b)]. Additional band structures for each material with an effective Hubbard are provided in the Supplemental Material [supplement]. The calculated bands are shown with colors indicating weight from uranium/lutetium (teal), chromium (red), or germanium (pink) states. The general features of the UCr6Ge6 band structure are, unsurprisingly, similar to those reported for other 166s [thomas2025uv6sn6, riedel2025magnetic, hu2022tunable, ortiz2024stability, kim2023infrared], but for UCr6Ge6, flatbands typically associated with the kagome lattice along the – path sit below . The relative shift also places a flat valence band with hybrid uranium and chromium character along the – Brillouin zone path close to , further demonstrating the band tuning possibilities provided by expanding the 166 family to 5 systems.
Comparing the UCr6Ge6, LuCr6Ge6, and previously published YCr6Ge6 [ishii2013ycr6ge6] band structures reveals that the Cr6Ge6 structures are conducive to a significant DOS at . The total DOS at for a UCr6Ge6 formula unit is 21.64 states/eV (13.79 states/eV for LuCr6Ge6). The equation
relates the Sommerfeld coefficient to the DOS at [] and the Boltzmann constant () for noninteracting electrons. The predicted Sommerfeld coefficient is 51.0 mJ mol-1 K-2 for UCr6Ge6 and 32.5 mJ mol-1 K-2 for LuCr6Ge6. The ratio of the experimental to predicted value is, therefore, 1.7 for UCr6Ge6 and 1.8 for LuCr6Ge6, indicating moderate enhancement. While the enhancement may indicate mass renormalization due to electron correlations, the increase may also stem from electron-phonon coupling. When an effective Hubbard is included in the band structure calculation to localize the uranium 5 electrons, the chromium DOS near is mostly unchanged, and the uranium weight at decreases to nearly zero (Fig. S5(a) [supplement]), resulting in a predicted Sommerfeld coefficient of 26.4 mJ mol-1 K-2.
III.4 Resistivity
Figure 7(a) shows the zero-field longitudinal resistivity of UCr6Ge6 with current applied in the kagome plane () and perpendicular to it (). The residual resistivity ratio [(300 K)/(1.8 K)] is 1.6 for and 3.7 for . Both are lower than the reported ratios for other Cr6Ge6 crystals, the structures of which are less disordered [konyk2020electrical, romaka2024structure], but the values are similar to those of UNb6Sn6 (3.3 [riedel2025magnetic]) and UV6Sn6 (2-3 [patino2025incom, thomas2025uv6sn6]). Neither orientation’s resistivity contains a feature near 30 K, and the first derivative of both orientations only shows a broad hump at the inflection point of the curve (Fig. S6 [supplement]). The resistivity with current along the axis () is lower than the resistivity with current within the plane (), as is also observed for YCr6Ge6 [ishii2013ycr6ge6] and for several CoSn-structure compounds [meier2020flat]. Figure 7(b) shows the magnetoresistance (MR) up to 9 T for both current directions with a perpendicular applied magnetic field. The magnetoresistance was calculated using
As with the zero-field, temperature-dependent resistivity, no transitions appear. The changes are small and monotonic for both orientations and are consistent with preferred conduction perpendicular to the kagome lattice.
IV Conclusions
UCr6Ge6 crystallizes in a structure with an average Y0.5Co3Ge3-type unit cell and a modulation that lowers the space-group symmetry to monoclinic. The modulation [] of UCr6Ge6 is distinct from other 166 materials in that it contains an component and no component. The isotropic temperature and magnetic field dependence of the magnetization contrast with other Cr6Ge6 and uranium-containing 166 materials with localized magnetic moments. Instead, the shape of the high-temperature susceptibility and the small magnetization strongly suggest that UCr6Ge6 is Pauli paramagnetic, though a ground state singlet due to a CEF splitting of a U 5 multiplet cannot be ruled out. A transition at 30 K in the magnetic susceptibility and heat capacity are attributed to an impurity phase. The electronic heat capacity, quantified by a Sommerfeld coefficient of 86.5 mJ mol-1 K-2, is moderately enhanced in UCr6Ge6 possibly due to flatbands near the Fermi level or electron-phonon coupling. Electronic band structure calculations suggest that the kagome flatbands and uranium states may increase the DOS near . The anisotropic resistivity and magnetoresistance reveal a preference for conduction perpendicular to the kagome lattice, consistent with other CoSn-type and stuffed-CoSn-type materials. The moderately enhanced Sommerfeld coefficient, isotropic magnetic properties, and unique structural modulation distinguish UCr6Ge6 from other uranium-containing 166 compounds and demonstrate the promise of tuning electronic band structure features and magnetic ground states in 5 166 materials.